The area of a circle is 100 times the area of another circle. What is the ratio of their circumferences?

Answer by thair(3)

asume c1 is the circumstans of small circle c2 is the circumstans for the big one
and a1 ,a2 are area for small and big circle
we have a1=(c1/2)r1
a2=100a1
then
(c2/2)*r2=100(c1/2)r1
c2/c1=100r1/r2 where r1 and r2 are the reduis of circles small and big